OpenStudy (anonymous):

10*9*8*7*6*5*4*3*2*1 = 10! Can this be true?! Why or why not?

OpenStudy (solomonzelman):

THis is true!

OpenStudy (solomonzelman):

That's the meaning of the "!" \[4!= 1\times 2 \times 3 \times 4~~~~~~~~or\]\[6!=1 \times 2 \times 3 \times 4 \times 5 \times 6\] See?

OpenStudy (anonymous):

Genius.

OpenStudy (solomonzelman):

That's just the definiton; ty!!